The owner of Bubba’s Barbeque is trying to determine a standard time for assembling each kind of sandwiches he sells. During a time study for assembling Bubba’s Famous Pulled Pork Sandwiches you recorded the following data in seconds.

Work Element Cycles Performance Rating
1 2 3 4 5
Prepare hoagie roll, move to prep area, open roll. 31 26 33 34 31 105%
Get pork from smoker. 26 19 25 26 19 95%
Shred pork for sandwich 149 148 145 140 141 100%
Place pork on roll. 6 5 6 6 7 103%
Pour on special BBQ sauce. 10 12 11 12 10 110%
Close hoagie roll, slice in half. 15 17 16 18 14 105%
Wrap sandwich, place in pickup area. 20 19 21 22 18 108%
Allowances per 8-hour shift
Change into uniform 11 minutes
Unavoidable delays 3 minutes
Break 10 minutes
Lunch 25 minutes
Change out of uniform 11 minutes
Total 60 minutes

a.   Compute the normal time for each element (Nt). Based on that information, what is the job's NT.  (Round your answer to two decimals.)


b.   Compute the allowance factor (for the workday/time worked) based on the allowance information provided.   % (Round your answer to one decimal. Do not include the % sign on your answer.)


c.   Compute standard time for the job.   (Round your answer to two decimals.)


d.   If the owner expects demand for Bubba’s Famous Pulled Pork Sandwiches to be 105 per hour, how many employees should be assigned to prepare this kind of sandwich?  
Answer #1

a.

B C D E F 1 31 2 26 33 25 4 34 26 19 A 1 Work element 2 Prepare hoagie roll, move to prep area, open roll 3 Get pork from smo

Formula

b.

Allowance factor = (60/480)*100 = 12.5%

c.

Standard time = total normal time/(1-allowance factor)

=255.68/(1-12.5%) = 292.2057143 = 292.21 seconds (Rounded to 2 decimal places)

d.No of employees required = 105/(3600/292.2057143) = 8.522666667 = 9 (Rounded to nearest whole number)